OpenStudy (anonymous):

Solve the equation. Check for extraneous solutions. [(x-4)/(x-1)]=[(10)/(x-7)] Please show work b/c i don't understand how to solve this correctly. Thank you.

OpenStudy (anonymous):

multiply, combine like terms , and set the equation equal to 0

OpenStudy (anonymous):

cross multiply?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{x-4}{x-1}=\frac{10}{x-7}\]Multiply both sides by (x-7) and then by (x-1)\[(x-7)(x-4) = 10(x-1)\]\[x^{2}-11x+28 = 10x-10\]\[x^{2}-21x+38 = 0\]Factor the quadratic:\[(x-19)(x-2)=0\]So x is 19 or 2

OpenStudy (anonymous):

factor and .... too late the solution is up ...

OpenStudy (anonymous):

check solutions to make sure they are not extraneous ...